Could you please help me with this. (1.10^x=12). I need to solve for x.
Answer by AnlytcPhil(1810)

1.10x = 12

Take the log of both sides

log(1.10x) = log(12) 

Use this rule of logarithms: log(AB) = B·log(A)

on the left side:

x·log(1.10) = log(12)

Divide both sides by log(1.10)

x·log(1.10)    log(12)
——————————— = —————————
 log(1.10)    log(1.10)

Cancel away the logs on the left:

      1
x·log(1.10)    log(12)
——————————— = —————————
 log(1.10)    log(1.10) 
    1


               log(12)
          x = —————————
              log(1.10)

Now get your calculator and get

               1.079181246
          x = ————————————— = 26.0717864
               .0413926852

Edwin
AnlytcPhil@aol.com